OpenStudy (anonymous):

Determine whether the series is convergent or divergent. If it is convergent, find its sum.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (1+3^n)/2^n\]

OpenStudy (anonymous):

no chance

OpenStudy (anonymous):

1/2^n + (3/2)^n

OpenStudy (anonymous):

(3/2)^n should be a tip-off...

OpenStudy (anonymous):

\[\frac{1+3^n}{2^n}>\frac{3^n}{n^n}\] terns don't even go to zero

OpenStudy (anonymous):

So what does that mean?

OpenStudy (anonymous):

typo there, but you get the idea \[\frac{1+3^n}{2^n}>\frac{3^n}{2^n}\] *terms don't go to zero

OpenStudy (anonymous):

a necessary (but not sufficient) condition for the sum to converge is that the terms go to zero

OpenStudy (anonymous):

it's a geometric series with |r| >1

OpenStudy (anonymous):

Okay, thanks.